# V-OCBF: Learning Safety Filters from Offline Data via Value-Guided Offline Control Barrier Functions

### v1

@@ Abstract
Ensuring safety in autonomous systems requires controllers that satisfy hard, state-wise constraints without relying on online interaction. While existing Safe Offline RL methods typically enforce soft expected-cost constraints, they do not guarantee forward invariance. Conversely, Control Barrier Functions (CBFs) provide rigorous safety guarantees but usually depend on expert-designed barrier functions or full knowledge of the system dynamics. We introduce Value-Guided Offline Control Barrier Functions (V-OCBF), a framework that learns a neural CBF entirely from offline demonstrations. Unlike prior approaches, V-OCBF does not assume access to the dynamics model; instead, it derives a recursive finite-difference barrier update, enabling model-free learning of a barrier that propagates safety information over time. Moreover, V-OCBF incorporates an expectile-based objective that avoids querying the barrier on out-of-distribution actions and restricts updates to the dataset-supported action set. The learned barrier is then used with a Quadratic Program (QP) formulation to synthesize real-time safe control. Across multiple case studies, V-OCBF yields substantially fewer safety violations than baseline methods while maintaining strong task performance, highlighting its scalability for offline synthesis of safety-critical controllers without online interaction or hand-engineered barriers.
@@ 1 Introduction
Control-theoretic tools provide an alternative and more rigorous foundation for safety. In particular, Control Barrier Functions (CBFs) (ames2014control) offer a principled mechanism to enforce hard, instantaneous safety constraints by guaranteeing the forward invariance of a prescribed safe set. When combined with learning-based controllers, CBFs serve as minimally invasive safety filters that adjust nominal actions only when necessary to prevent constraint violations. Their integration with Quadratic Program (QP) based controllers enables real-time implementation with modern optimization solvers. Consequently, CBF-based controllers have been successfully applied to a wide range of safety-critical tasks, including adaptive cruise control (ames2014control), aerial robotics (7525253; tayal2024control), and legged locomotion (nguyen2015safety). In all of these applications, the performance and safety guarantees fundamentally depend on the quality of the underlying CBF.
@@
Recent work has explored learning Control Barrier Functions (CBFs) from offline demonstrations (Alexander2020NCBF; Fernando2023iDBF; tabbara2025CCBF). Existing methods either fit CBFs only on expert trajectories or rely on data-likelihood measures to filter unsafe samples, which limits their ability to generalize beyond the demonstrated states. Uncertainty-aware approaches address distributional mismatch but often become overly conservative. Overall, current offline CBF learning methods are closely tied to the empirical data distribution and do not explicitly reason about future system evolution, resulting in conservative safety guarantees.
This paper introduces Value-Guided Offline Control Barrier Functions (V-OCBF), a novel framework designed to overcome key limitations of existing offline RL and CBF-based approaches. We derive a model-free finite-difference recursion for updating the barrier function, and we show that satisfying this update provides a formal one-step safety guarantee for any control-affine system under the resulting policy. In addition, we propose an expectile-based learning objective that allows the synthesized safe policy to improve over the behavior policy in the dataset while never querying the barrier on out-of-distribution actions, ensuring stable and reliable offline learning. To summarise, the main contributions of this work are as follows:
@@
    We propose V-OCBF, a framework for learning formally safe controllers entirely from offline demonstrations.
@@ 2.3 Positioning of Our Work
V-OCBF differs from safe offline RL methods (lee2022coptidice; liu2023constrained; lin2023safe; 1), which rely on soft constraints, by directly constructing a hard safety certificate that guarantees one-step forward invariance. Unlike offline neural CBF approaches (Alexander2020NCBF; Fernando2023iDBF; tabbara2025CCBF) that remain tied to the demonstration distribution, V-OCBF employs a model-free value recursion that models future safety evolution and expands the safe set beyond observed states. Overall, our method combines the distributional robustness of offline RL with the rigor of CBF-based control to obtain an offline-learned, actuation-aware neural CBF without requiring online rollouts.
@@ 3 Background and Problem Setup
Since, $\mathcal{F}\subseteq\mathcal{X}\setminus\mathcal{C}$, it can be trivially shown that $x(t)\in\mathcal{C}\implies x(t)\notin\mathcal{F}~\forall~t\geq 0$. Using this premise, we define the main objective of this paper: {objective} Our objective is to synthesize a safe policy $\pi_{\mathrm{safe}}:[t,T)\times\mathcal{X}\to\mathbb{U}$ such that the resulting closed-loop system satisfies the positive invariance property specified in Definition equation 1.
@@ 3.2 Control Barrier Value Function (CBVF)
Our objective is to synthesize a Control Barrier Value Function $B:\mathcal{X}\to\mathbb{R}$ directly from an offline dataset of demonstrations $\mathcal{D}$, such that the resulting CBVF-QP controller $\pi_{\mathrm{safe}}$ in equation 7 satisfies the positive invariance property specified in Definition 1.
@@ 4.1 Finite-Difference Barrier Synthesis
where, $x_{t}=x(t)$, $u_{t}=u(t)$ and $x_{t+1}=x(t+\Delta t)$, along with the boundary condition $B(x_{0})=\ell(x_{0})$. This recursion has two significant advantages. First, it enables us to learn a barrier directly from data without requiring $f$ and $g$. Second, under mild regularity assumptions, equation 8 preserves the forward invariance property of the CBVF. The derivation of the recursive equation is provided in Appendix A.1. Intuitively, the recursion encodes the principle that a state is safe if and only if its immediate successor is safe or it lies outside the unsafe set as specified by $\ell(x)$.
To represent this barrier function, we parameterize a neural network $B^{\mathcal{D}}_{\psi}(x)$ with $\psi$ utilizing the universal approximation property. However, directly solving for the recursion equation 8 can lead to degenerate solutions. For instance, setting $B^{\mathcal{D}}_{\psi}(x)=c$ for a sufficiently small constant $c$ satisfies equation 8 but clearly does not satisfy the CBVF conditions in equation 5. This pathology is analogous to the non-contractive behavior of undiscounted value iteration in MDPs. Following the approach in (Fisac2019HJSafety), we incorporate a discounted finite-difference loss to avoid such trivial solutions:
$\mathcal{L}_{B^{\mathcal{D}}}(\psi)=\mathbb{E}_{(x,u,x^{\prime})\sim\mathcal{D}}\big[\big((1-\gamma)\,\ell(x)+\gamma\min\big\{\ell(x),\,B^{\mathcal{D}}_{\psi}(x^{\prime})\big\}-B^{\mathcal{D}}_{\psi}(x)\big)^{2}\big],$ (9)
where $x_{t}, x=x$$^{\prime}=x_{t+1}$ and $u=u_{t}$ and $\gamma\rightarrow 1$. This discounted recursion ensures contraction, promotes stable learning, and prevents the network from collapsing to uniformly unsafe or uniformly safe solutions. To avoid evaluating barrier targets at out-of-distribution actions, we remove the maximization over actions from the loss in equation 9 and use only the demonstrated action in each transition. While this prevents unsupported queries, the resulting estimate reflects the safety profile of the behaviour policy that generated the data. Consequently, this produces a behaviour-induced barrier, which is typically sub-optimal because it ignores other admissible actions that could yield larger safe-set estimates.
@@ 4.2 Avoiding Out-of-Distribution Actions in Offline Learning
The naïve regression objective in equation 9 fits $B^{\mathcal{D}}_{\psi}$ to the mean of the demonstrated next-state targets, but this corresponds to the behavior-induced barrier and yields overly conservative safe sets. Ideally, if we assume unlimited capacity and no sampling error, the optimal parameters should satisfy, $B(x)\approx\mathbb{E}_{u}\big[\min\big\{\ell(x),\,\max_{u}B(x^{\prime})\big\}\big]$. However, such unconstrained maximization can result in actions that are never observed in the dataset.
Since offline data only provides information about those actions selected by the behavior policy, evaluating values (or barrier targets) using unsupported actions can distort learning because the corresponding transitions are not grounded in the dataset. Subsequently, motivated by the insights of Implicit Q-Learning (IQL) (ilya2022IQL), we approximate the maximization over admissible actions by using expectile regression, which enables us to capture the highest admissible barrier values supported by the data while never evaluating $B$ on unseen $(x,u)$ pairs. This allows us to perform a principled value-style backup over the dataset control action support $\mathcal{U}_{\mathcal{D}}=\;\{\,u\mid(x,u,x^{\prime})\in\mathcal{D}\,\}$ without extrapolating to unsafe or unobserved actions. IQL shows that expectile regression produces a value function that reflects the values induced by the behavior policy without requiring an explicit behavior model. This prevents the learning target from being influenced by actions that lie outside the dataset support, while still capturing the highest feasible values supported by the demonstrations.
@@
$\mathcal{L}_{B}(\theta)\;=\;\mathbb{E}_{(x,u,x^{\prime})\sim\mathcal{D}}\big[\mathcal{L}^{\tau}\!\big(B^{\mathcal{D}}_{\psi}(x)-B_{\theta}(x)\big)\big],$ (10)
where $\mathcal{L}^{\tau}(y)=\lvert\tau-\mathbf{1}(y<0)\rvert\,y^{2}$ is the $\tau$-expectile loss used in (ilya2022IQL). Intuitively, a higher expectile level $\tau$ places greater weight on underestimation errors than overestimation errors, pushing $B_{\theta}$ toward the upper envelope of safety values supported by the dataset. Thus, $\tau$ controls how aggressively the learned barrier emphasizes high, data-supported safety values without extrapolating to unseen actions. The barrier function thus obtained, $B_{\theta}$, is our proposed Value-guided Offline Control Barrier Function (V-OCBF). The complete learning procedure for $B^{\mathcal{D}}_{\psi}$ and the V-OCBF $B_{\theta}$ is summarized in Algorithm 1 and can be visually seen in Fig. 1.
@@
5:   (A) $B^{\mathcal{D}}_{\psi}$ update
6:   $\mathcal{L}_{B^{\mathcal{D}}}\leftarrow$$(x,u,x^{\prime},\gamma,\psi)$ with $(x,u,x^{\prime})\sim\mathcal{B}$ $\triangleright$ From Eq. 9
7:   $\psi\leftarrow\text{AdamStep}(\psi,\nabla_{\psi}\,\mathcal{L}_{B^{\mathcal{D}}})$
8:   (B) form support targets
9:   For each unique $x\in\mathcal{D}$ gather support $\mathcal{U}_{\mathcal{D}}$ from $\mathcal{D}$
10:   (C) B-network losses
11:   $\mathcal{L}_{B}\leftarrow$ $(x,\mathcal{U}_{\mathcal{D}},\tau,B^{\mathcal{D}}_{\psi},\theta)$ $\triangleright$ From Eq. 10
12:   $\theta\leftarrow\text{AdamStep}(\theta,\nabla_{\theta}\mathcal{L}_{B})$
13:  end for
14:end for
15:Return $B_{\theta}$
@@ 4.3 Controller Synthesis via Learned Dynamics
These quantities allow the QP in equation 7 to be solved for the safe control action $u_{safe}$, completing the pipeline for constructing a safety-certified controller purely from offline demonstrations.
In Section 5.2.2, we provide an analysis demonstrating the limitations of using learned dynamics inside the CBVF learning loop, further reinforcing the necessity of restricting their use to the inference stage only.
@@ 5 Experiments
Baselines: We compare V-OCBF against a diverse set of constrained offline learning and CBF-based methods. For constrained offline learning, we include Behavior Cloning (BC), BEAR-Lag (Lagrangian constraint version of (kumar2019stabilizing)), COptiDICE (lee2022coptidice), and FISOR (1) which enforce safety indirectly via soft constraints on policy optimization or behavior imitation. For CBF-based approaches, we evaluate Neural Control Barrier Function (NCBF) (Alexander2020NCBF), Conservative Control Barrier Function (CCBF) (tabbara2025CCBF), and In-Distribution Barrier Function (iDBF) (Fernando2023iDBF), which synthesize explicit safety filters from offline data but often yield conservative safe sets. In contrast, V-OCBF learns a value-guided barrier function from offline demonstrations that accounts for future unsafe interactions, producing a hard, state-wise safety filter with larger safe set coverage.
@@ 5.2.2 Effect of learned dynamics
5.3 Ablation Studies
To further evaluate the robustness of V-OCBF, we conduct ablation studies that examine: (i) the sensitivity of the learned safe set to the size of the barrier network, (ii) the effect of disturbances on the safety of QP based controller, and (iii) the dependence of the learned barrier on the expectile parameter $\tau$. Across these experiments, V-OCBF exhibits stable safety performance over a broad range of architectural and hyperparameter choices, remaining resilient even under perturbations to the control actions. These findings suggest that the framework generalises reliably beyond the specific tasks considered and is well-suited for settings where safety certification must be carried out under imperfect system knowledge. The full results and visualisations for these studies are presented in Appendix C.

### v2

@@ Abstract
Ensuring safety in autonomous systems requires controllers that aim to satisfy state-wise constraints without relying on online interaction.While existing Safe Offline RL methods typically enforce soft expected-cost constraints, they struggle to ensure strict state-wise safety. Conversely, Control Barrier Functions (CBFs) offer a principled mechanism to enforce forward invariance, but often rely on expert-designed barrier functions or knowledge of the system dynamics. We introduce Value-Guided Offline Control Barrier Functions (V-OCBF), a framework that learns a neural CBF entirely from offline demonstrations. Unlike prior approaches, V-OCBF does not assume access to the dynamics model; instead, it derives a recursive finite-difference barrier update, enabling model-free learning of a barrier that propagates safety information over time. Moreover, V-OCBF incorporates an expectile-based objective that avoids querying the barrier on out-of-distribution actions and restricts updates to the dataset-supported action set. The learned barrier is then used with a Quadratic Program (QP) formulation to synthesize real-time safe control. Across multiple case studies, V-OCBF yields substantially fewer safety violations than baseline methods while maintaining strong task performance, highlighting its scalability for offline synthesis of safety-critical controllers without online interaction or hand-engineered barriers.
@@ 1 Introduction
Control-theoretic tools provide an alternative and more rigorous foundation for safety. In particular, Control Barrier Functions (CBFs) (ames2014control) offer a principled mechanism to enforce instantaneous safety constraints by guaranteeing the forward invariance of a prescribed safe set. When combined with learning-based controllers, CBFs serve as minimally invasive safety filters that adjust nominal actions only when necessary to prevent constraint violations. Their integration with Quadratic Program (QP) based controllers enables real-time implementation with modern optimization solvers. Consequently, CBF-based controllers have been successfully applied to a wide range of safety-critical tasks, including adaptive cruise control (ames2014control), aerial robotics (7525253; tayal2024control), and legged locomotion (nguyen2015safety). In all of these applications, the performance and safety guarantees fundamentally depend on the quality of the underlying CBF.
@@
Recent work has explored learning Control Barrier Functions (CBFs) from offline demonstrations (Alexander2020NCBF; Fernando2023iDBF; tabbara2025CCBF). Existing methods either fit CBFs only on expert trajectories or rely on data-likelihood measures to filter unsafe samples, which limits their ability to generalize beyond the demonstrated states. Uncertainty-aware approaches address distributional mismatch but often become overly conservative. Overall, current offline CBF learning methods are closely tied to the empirical data distribution and do not explicitly reason about future system evolution, resulting in conservative safety.
This paper introduces Value-Guided Offline Control Barrier Functions (V-OCBF), a novel framework designed to overcome key limitations of existing offline RL and CBF-based approaches. We derive a model-free finite-difference barrier recursion and establish that, in the idealized setting without learning or approximation errors, enforcing this update provides a formal one-step forward-invariance safety guarantee for control-affine systems. In addition, we propose an expectile-based learning objective that allows the synthesized safe policy to improve over the behavior policy in the dataset while never querying the barrier on out-of-distribution actions, ensuring stable and reliable offline learning. To summarise, the main contributions of this work are as follows:
@@
    We propose V-OCBF, a framework for learning safe controllers entirely from offline demonstrations.
@@ 2.3 Positioning of Our Work
V-OCBF positions itself primarily as a framework for learning safe control policies from offline datasets, utilizing Control Barrier Functions (CBFs) as a learned intermediate representation to enforce strictly safe behavior. Unlike safe offline RL methods (lee2022coptidice; liu2023constrained; lin2023safe; 1), which rely on soft constraints, we directly construct a safety certificate and a safe policy that guarantees one-step forward invariance, in the idealised setting. In contrast to model-based safe offline approaches that rely on learned dynamics both for training neural CBFs and during inference (Alexander2020NCBF; Fernando2023iDBF; tabbara2025CCBF), V-OCBF employs a model-free value recursion that models future safety evolution and uses dynamics only at the time of inference for local Lie derivative computation in the CBF-QP, rather than for policy optimization or imagined rollouts. This design prevents the policy from extrapolating into action regions unsupported by the dataset and preserves the OOD robustness of our expectile-based objective. Overall, our method combines the distributional robustness of offline RL with the rigor of CBF-based control to obtain an offline-learned, actuation-aware neural CBF without requiring online rollouts.
@@ 3 Background and Problem Setup
Since, $\mathcal{F}\subseteq\mathcal{X}\setminus\mathcal{C}$, it can be trivially shown that $x(t)\in\mathcal{C}\implies x(t)\notin\mathcal{F}~\forall~t\geq 0$. Using this premise, we define the main objective of this paper: {objective} Our objective is to synthesize a safe policy $\pi_{\mathrm{safe}}:[t,T)\times\mathcal{X}\to\mathbb{U}$ such that the resulting closed-loop system satisfies the positive invariance property specified in Definition 1.
@@ 3.2 Control Barrier Value Function (CBVF)
Our objective is to synthesize a Control Barrier Value Function $B:\mathcal{X}\to\mathbb{R}$ directly from an offline dataset of demonstrations $\mathcal{D}$, such that the resulting CBVF-QP controller $\pi_{\mathrm{safe}}$ in equation 7 aims to satisfy the positive invariance property specified in Definition 1.
@@ 4.1 Finite-Difference Barrier Synthesis
where, $x_{t}=x(t)$, $u_{t}=u(t)$ and $x_{t+1}=x(t+\Delta t)$. This recursion has two significant advantages. First, it enables us to learn a barrier directly from data without requiring $f$ and $g$. Second, under mild regularity assumptions, equation 8 preserves the forward invariance property of the CBVF. The derivation of the recursive equation is provided in Appendix A.1. Intuitively, the recursion encodes the principle that a state is safe if and only if its immediate successor is safe or it lies outside the unsafe set as specified by $\ell(x)$.
To represent this barrier function, we parameterize a neural network $B^{\mathcal{D}}_{\psi}(x)$ with $\psi$ utilizing the universal approximation property. However, directly solving for the recursion equation 8 can lead to degenerate solutions. For instance, setting $B^{\mathcal{D}}_{\psi}(x)=c$ for a sufficiently small constant $c$ satisfies equation 8 but clearly does not satisfy the CBVF conditions in equation 5. This pathology is analogous to the non-contractive behavior of undiscounted value iteration in MDPs. Following the approach in (Fisac2019HJSafety), we incorporate a discounted finite-difference value to avoid such trivial solutions:
$B^{\mathcal{D}}_{Target}=(1-\gamma)\,\ell(x)+\gamma\min\big\{\ell(x),\,\max_{u}B^{\mathcal{D}}_{\psi}(x^{\prime})\big\},$ (9)
where $x_{t}, x=x$$^{\prime}=x_{t+1}$ and $u=u_{t}$ and $\gamma\rightarrow 1$. This discounted recursion ensures contraction, promotes stable learning, and prevents the network from collapsing to uniformly unsafe or uniformly safe solutions. To avoid evaluating barrier targets at out-of-distribution actions, we remove the maximization over actions from the equation 9 and use only the demonstrated action in each transition. While this prevents unsupported queries, the resulting estimate reflects the safety profile of the behaviour policy that generated the data. Consequently, this produces a behaviour-induced barrier, which is typically sub-optimal because it ignores other admissible actions that could yield larger safe-set estimates.
@@ 4.2 Avoiding Out-of-Distribution Actions in Offline Learning
The naïve regression objective in equation 9 with the maximization over actions dropped, fits $B^{\mathcal{D}}_{\psi}$ to the mean of the demonstrated next-state targets, but this corresponds to the behavior-induced barrier and yields overly conservative safe sets. Ideally, if we assume unlimited capacity and no sampling error, the optimal parameters should satisfy, $B(x)\approx\mathbb{E}_{u}\big[\min\big\{\ell(x),\,\max_{u}B(x^{\prime})\big\}\big]$. However, such unconstrained maximization can result in actions that are never observed in the dataset.
Since offline data only provides information about those actions selected by the behavior policy, evaluating values (or barrier targets) using unsupported actions can distort learning because the corresponding transitions are not grounded in the dataset. Subsequently, motivated by the insights of Implicit Q-Learning (IQL) (ilya2022IQL), under idealized conditions, deterministic dynamics and disturbance-free data, prioritizing the best observed next-state outcomes can yield strong empirical performance, since emphasizing high-value targets leads to safer barrier estimates. However, this naïve maximization is brittle in practice: rare poor actions or stochastic disturbances can occasionally produce optimistic next states, and focusing only on the single best outcome can cause the learner to overfit such spurious transitions. Expectile regression provides a simple remedy by emphasizing high next-state barrier values that actually appear under dataset-supported actions, rather than committing to a single extreme outcome. In this way, it preserves the benefit of prioritizing good actions while remaining robust to noise and avoiding evaluations on unseen $(x,u)$ pairs. For details and sensitivity to $\tau$, see Appendix A.2 and Appendix C.3, respectively. This allows us to perform a principled value-style backup over the dataset control action without extrapolating to unsafe or unobserved actions. This IQL-inspired method shows that expectile regression produces a value function that reflects the values induced by the behavior policy without requiring an explicit behavior model. This prevents the learning target from being influenced by actions that lie outside the dataset support, while still capturing the highest feasible values supported by the demonstrations.
@@
$\mathcal{L}_{B}(\theta)\;=\;\mathbb{E}_{(x,u,x^{\prime})\sim\mathcal{D}}\big[\mathcal{L}^{\tau}\!\big(B^{\mathcal{D}}_{Target}(x)-B_{\theta}(x)\big)\big],$ (10)
where $\mathcal{L}^{\tau}(y)=\lvert\tau-\mathbf{1}(y<0)\rvert\,y^{2}$ is the $\tau$-expectile loss used in (ilya2022IQL) and we replace the original objective term $B^{\mathcal{D}}_{Target}$ from equation 9 with the new objective $B^{\mathcal{D}}_{Target}=(1-\gamma)\ell(x)+\gamma\min\{\ell(x),B_{\theta}(x^{\prime})\}$ which incorporates $B_{\theta}(x^{\prime})$ in place of $\max_{u}B^{\mathcal{D}}_{\psi}(x^{\prime})$. Intuitively, a higher expectile level $\tau$ places greater weight on underestimation errors than overestimation errors, pushing $B_{\theta}$ toward the upper envelope of safety values supported by the dataset. Thus, $\tau$ controls how aggressively the learned barrier emphasizes high, data-supported safety values without extrapolating to unseen actions. In practice, this is implemented by computing the expectile loss over the empirical state-action transitions within each sampled minibatch, thereby restricting the updates to the dataset-supported action set without requiring explicit support construction. The barrier function thus obtained, $B_{\theta}$, is our proposed Value-guided Offline Control Barrier Function (V-OCBF). The complete learning procedure for $B^{\mathcal{D}}_{\psi}$ and the V-OCBF $B_{\theta}$ is summarized in Algorithm 1 and can be visually seen in Fig. 1.
@@
5:   $\mathcal{L}_{B}\leftarrow\frac{1}{|\mathcal{B}|}\sum_{(x,u,x^{\prime})\in\mathcal{B}}\mathcal{L}^{\tau}\left(B^{\mathcal{D}}_{Target}(x)-B_{\theta}(x)\right)$ $\triangleright$ From Eq. 10
6:   $\theta\leftarrow\text{AdamStep}(\theta,\nabla_{\theta}\mathcal{L}_{B})$
7:  end for
8:end for
9:Return $B_{\theta}$
@@ 4.3 Controller Synthesis via Learned Dynamics
These quantities allow the QP in equation 7 to be solved for the safe control action $u_{safe}$, completing the pipeline for constructing a safe controller purely from offline demonstrations.
@@
Note that one-step forward-invariance holds under the idealized assumptions (exact barrier and exact Lie-derivatives), and the approximation sources like function approximation, finite-data estimation, and model error can break the closed-loop guarantee in practice. In Section 5.2.2, we provide an analysis demonstrating the limitations of using learned dynamics inside the CBVF learning loop, further reinforcing the necessity of restricting their use to the inference stage only.
@@ 5 Experiments
Baselines: We compare V-OCBF against a diverse set of constrained offline learning and CBF-based methods. For constrained offline learning, we include Behavior Cloning (BC), BEAR-Lag (Lagrangian constraint version of (kumar2019stabilizing)), COptiDICE (lee2022coptidice), and FISOR (1) which enforce safety indirectly via soft constraints on policy optimization or behavior imitation. For CBF-based approaches, we evaluate Neural Control Barrier Function (NCBF) (Alexander2020NCBF), Conservative Control Barrier Function (CCBF) (tabbara2025CCBF), and In-Distribution Barrier Function (iDBF) (Fernando2023iDBF), which synthesize explicit safety filters from offline data but often yield conservative safe sets. In contrast, V-OCBF learns a value-guided barrier function from offline demonstrations that accounts for future unsafe interactions, producing a state-wise safety filter that aims to reduce violations and increase safe-set coverage in practice.
@@ 5.2.2 Effect of learned dynamics
5.3 Ablation Studies and Additional Experiments
To further evaluate the robustness of V-OCBF, we conduct a series of studies that examine: (i) the sensitivity of the learned safe set to the size of the barrier network, (ii) the effect of disturbances on the safety of QP based controller, (iii) the dependence of the learned barrier on the expectile parameter $\tau$, and (iv) the resilience of the framework to approximation errors in the learned dynamics model. Additionally, we extend our evaluation to a boat navigation task with nonlinear drift to assess performance under complex environmental dynamics. Across these experiments, V-OCBF exhibits stable safety performance over a broad range of architectural and hyperparameter choices, remaining resilient even under perturbations to the control actions. These findings suggest that the framework generalises reliably beyond the specific tasks considered and is well-suited for settings where safety certification must be carried out under imperfect system knowledge. The full results and visualisations for these studies are presented in Appendix C and Appendix D.